solve the system of equation x+y+z=9
-x+y+z=1
x-y-z=5
would appreciate some help

Answer by Alan3354(69443)

solve the system of equation
x+y+z=9
-x+y+z=1
x-y-z=5
================
Adding eqns 2 & 3 --> 0 = 6